Solution A is 30% alcohol and solution B is 75% alcohol.How much of each should be used to make 100 liters of a solution that is 50% alcohol.
Answer by checkley77(12844)

.75x+.30(100-x)=.50 *100
.75x+30-.30x=50
.45x=50-30
.45x=20
x=20/.45
x=44.444 liters of 75% alcohol is used.
100-44.444=55.556 liters of 30% alcohol is used.
Proof:
.75*44.444+.30*55.556=.50*100
33.333+16.667=50
50=50